Find the x-intercept of the graph of 4x - 2y = 8.
Answer by Theo(13342)

the x intercept is the value of x when y = 0.

your equation is:

4x - 2y = 8

replace y with 0 to get:

4x - 2 * 0 = 8

simplify to get:

4x = 8

solve for x to get:

x = 8/4 = 2

the x-intercept is equal to 2.

the coordinate point where this occurs is (x,y) = (2,0).

here's a graph of the equation.



you can see that x = 2 when y = 0

some additional information.

the y-intercept is the value of y when x = 0.

when x = 0, 4x - 2y = 8 becomes 4 * 0 - 2y = 8.

simplify to get:

-2y = 8

solve for y to get:

y = 8 / -2 = -4

the coordinate point where this occurs is (x,y) = (0,-4)

both the x and y intercepts can be seen on the graph.